OpenStudy (anonymous):

Find the power series for the following and determine the radius of convergence. f(x) = x^2/(2+x) AWARD GIVEN AT END!

OpenStudy (zarkon):

\[\frac{1}{2+x}=\frac{1}{2\left(1+\dfrac{x}{2}\right)}=\frac{1}{2\left(1-\dfrac{-x}{2}\right)}\]

OpenStudy (zarkon):

\[\sum_{n=0}^{\infty}r^n=\frac{1}{1-r}\text{ for }|r|<1\]

OpenStudy (anonymous):

So f'x is 1/(2+x)? @Zarkon

OpenStudy (zarkon):

by the above we have ... \[\frac{1}{2\left(1-\dfrac{-x}{2}\right)}=\frac{1}{2}\sum_{n=0}^{\infty}(-x/2)^n\] for |-x/2|<1

OpenStudy (zarkon):

then multiply by \(x^2\) to get the full series

OpenStudy (anonymous):

\[\frac{1}{2}\sum_{n=0}^{\infty}(-x/2)^n \times x^2\] ??

OpenStudy (zarkon):

yes....then ... \[\Large\frac{1}{2}\sum_{n=0}^{\infty}(-x/2)^nx^2=\frac{1}{2}\sum_{n=0}^{\infty}(-x)^nx^2/2^n=\sum_{n=0}^{\infty}(-x)^{n+2}/2^{n+1}\]

OpenStudy (zarkon):

\[\Large=\sum_{n=0}^{\infty}(-1)^{n+2}x^{n+2}/2^{n+1}\] \[\Large=\sum_{n=0}^{\infty}\frac{(-1)^{n}x^{n+2}}{2^{n+1}}\]

OpenStudy (anonymous):

Awesome! :) Thanks! So the radius of convergence is 1?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

we need |-x/2|<1 so |x|<2 so the radius of convergence is 2

OpenStudy (anonymous):

Ohhh okay. Thanks a bunch! :)